# Scoring Time Intervals using Non-Hierarchical Transformer for Automatic Piano Transcription

### v5

@@ Abstract
The neural semi-Markov Conditional Random Field (semi-CRF) framework has demonstrated promise for event-based piano transcription. In this framework, all events (notes or pedals) are represented as closed time intervals tied to specific event types. The neural semi-CRF approach requires an interval scoring matrix that assigns a score for every candidate interval. However, designing an efficient and expressive architecture for scoring intervals is not trivial. This paper introduces a simple method for scoring intervals using scaled inner product operations that resemble how attention scoring is done in transformers. We show theoretically that, due to the special structure from encoding the non-overlapping intervals, under a mild condition, the inner product operations are expressive enough to represent an ideal scoring matrix that can yield the correct transcription result. We then demonstrate that an encoder-only structured non-hierarchical transformer backbone, operating only on a low-time-resolution feature map, is capable of transcribing piano notes and pedals with high accuracy and time precision. The experiment shows that our approach achieves the new state-of-the-art performance across all subtasks in terms of the F1 measure on the Maestro dataset.
@@ 3.1 Interval Scoring in [2]
After producing the interval scoring matrices for all event types, a shallow CNN is applied, treating the interval endpoints as spatial coordinates and event types as channels. This refinement step slightly improves the result.
@@ 3.3 Comparison with Attention Mechanism
Comparing the neural semi-CRF with the inner product scoring to the attention mechanism reveals interesting parallels. The original score module, as in [2], resembles an additive attention mechanism, as introduced by [12]. However, attention mechanisms based on inner products [13] have become preferred for their simplicity and computational efficiency. Similarly, the proposed inner product scoring for neural semi-CRFs efficiently scores intervals. Computationally, semi-CRFs have the same quadratic computational time complexity as attention heads. However, in contrast to attention mechanisms that score sequence positions and normalize posteriors for each position, neural semi-CRFs score intervals and normalize posteriors globally over sets of non-overlapping intervals.
@@ 5.1 Dataset
Similar to Maestro dataset, the SMD dataset was created by recording human performance on a Yamaha Disklavier. The dataset contains 50 recordings. We found that the onset annotations in SMD are better aligned compared to MAPS, but the offset annotations suffer from similar biases.
@@ 5.2 Model Specification
Strided Conv. Layers initial proj. size: 64, added with freq. embeddings
for Downsampling out: [128, 256, 256, 256], kernel size: 3, stride: [(2,1), (2,2), (2,2), (1,1)], Each followed by GroupNorm, groups = 4, and GELU (except for the last conv.)
@@ 5.3 Evaluation Metrics
For MAPS and SMD, we only report activation-level and onset-only note-level metrics due to the ground-truth alignment issues discussed in Section 5.1. These issues make offset-related metrics unreliable for these datasets.
@@ 5.4 Results
All methods exhibit low activation-level F1 scores on both MAPS and SMD. Using the onset-corrected annotation (Cogliati) on MAPS increases the onset F1 score but degrades the activation-level F1 score due to the uncorrected offset biases. In fact, the Cogliati annotation achieves similar or lower activation-level F1 scores compared to all listed methods when evaluated against the original annotation.
For note onset, all methods on SMD achieve F1 scores close to those evaluated on Maestro. However, performance decreases significantly on MAPS, even with corrected annotations, suggesting that the alignment issue may be more complex than a simple global offset.
@@
00footnotetext: Previous SOTA for notes. Use their provided code and pretrained weights. Recomputed from transcribed MIDIs. 00footnotetext: Previous SOTA for sustain pedals. Their released code indicates a 200 ms onset tolerance for pedal evaluation, contrary to the reported 50 ms in their paper. Here, we use a 50 ms onset tolerance, which explains the large discrepancy between the numbers here and their reported results.

### v6

@@ Abstract
The neural semi-Markov Conditional Random Field (semi-CRF) framework has demonstrated promise for event-based piano transcription. In this framework, all events (notes or pedals) are represented as closed time intervals tied to specific event types. The neural semi-CRF approach requires an interval scoring matrix that assigns a score for every candidate interval. However, designing an efficient and expressive architecture for scoring intervals is not trivial. This paper introduces a simple method for scoring intervals using scaled inner product operations that resemble how attention scoring is done in transformers. We show theoretically that, due to the special structure from encoding the non-overlapping intervals, under a mild condition, the inner product operations are expressive enough to represent an ideal scoring matrix that can yield the correct transcription result. We then demonstrate that an encoder-only non-hierarchical transformer backbone, operating only on a low-time-resolution feature map, is capable of transcribing piano notes and pedals with high accuracy and time precision. The experiment shows that our approach achieves the new state-of-the-art performance across all subtasks in terms of the F1 measure on the Maestro dataset.
See appendix for post-camera-ready updates.
@@ 3.1 Interval Scoring in [2]
After producing the initial interval scoring matrices for all event types, a shallow CNN is applied, treating the interval endpoints as spatial coordinates and event types as channels. This refinement step slightly improves the result.
@@ 3.3 Comparison with Attention Mechanism
Comparing the neural semi-CRF with the inner product scoring to the attention mechanism reveals interesting parallels. Both of them have quadratic time complexity in the length of the input. The original score module, as in [2], resembles an additive attention mechanism, as introduced by [12]. However, attention mechanisms based on inner products [13] have become preferred for their simplicity and computational efficiency. Similarly, the proposed inner product scoring for neural semi-CRFs efficiently scores intervals. However, in contrast to attention mechanisms that score sequence positions and normalize posteriors for each position, neural semi-CRFs score intervals and normalize posteriors globally over sets of non-overlapping intervals.
@@ 5.1 Dataset
Similar to Maestro dataset, the SMD dataset was created by recording human performance on a Yamaha Disklavier. We use SMD version 2. The dataset contains 50 recordings. We found that both the onset and offset annotations in SMD are better aligned compared to MAPS.
@@ 5.2 Model Specification
Strided Conv. Layers initial proj. size: 64, added with freq. embeddings.
for Downsampling out channels: [128, 256, 256, 256], kernel size: 3, strides: [(2,1), (2,2), (2,2), (1,1)], Each followed by GroupNorm, groups = 4, and GELU (except for the last conv.)
@@ 5.3 Evaluation Metrics
Due to the ground-truth alignment issues discussed in Section 5.1 and space constraints, we only report activation-level and onset-only note-level metrics for MAPS and SMD.
@@ 5.4 Results
All methods exhibit low activation-level F1 scores on MAPS. Using the onset-corrected annotation (Cogliati) on MAPS increases the onset F1 score but degrades the activation-level F1 score due to the uncorrected offset biases. In fact, the Cogliati annotation achieves similar or lower activation-level F1 scores compared to all listed methods when evaluated against the original annotation.
All methods achieve F1 scores on SMD that are more comparable to those evaluated on Maestro. However, performance decreases significantly on MAPS, even with corrected annotations. This suggests that the dataset issue may be more complex than a simple piece-depedent timing shift.
@@
00footnotetext: Use their provided code and pretrained weights. Recomputed from transcribed MIDIs. 00footnotetext: Previous SOTA for sustain pedals. Their released code indicates a 200 ms onset tolerance for pedal evaluation, contrary to the reported 50 ms in their paper. Here, we use a 50 ms onset tolerance, which explains the large discrepancy between the numbers here and their reported results.
@@ 7 Acknowledgement
Appendix A Changes from Previous Versions
We later discovered that the SMD version 1, which was used in the previous versions of the paper, missed all CC events, which affects all offsets of pedal-extended notes. We notified the original author and they provided the revised SMD version 2 [33]. We updated the numbers in our paper to reflect this fix.